Two years ago John was three times as old as Frank. Five years from now John will be twice as old as Frank. What is the sum of their ages right now?
how do you figure this out? thanks

Answer by Edwin McCravy(20054)

Two years ago John was three times as old as Frank. Five years from now John will be twice as old as Frank. What is the sum of their ages right now?
how do you figure this out? thanks
Two years ago John was

J-2, which was = to

three times

that's 3*

as old as Frank.

(was two years ago), which was (F-2).

So following everything in red above,
we have this equation:

J-2 = 3*(F-2)

Five years from now John will be

J+5, which =

twice as old

(that's 2*)

as Frank.

(as Frank will be five years from now), 
which will be (F+5) 

So following everything above in green, 
we have this equation:

J+5 = 2*(F+5)

So we have this system:

J-2 = 3*(F-2)
J+5 = 2*(F+5)

Can you simplify that system and 
solve it?  If not post again asking how.

J=11, F=5

So John is 11 and Frank is 5

Checking:

Two years ago John was

(9, which was) 

three times as old as Frank.

(who was then 3) 

Five years from now John will be

(16, which will be) 

twice as old as Frank.

(who will be 8).

What is the sum of their ages right now?

Add 11 + 5 and get 16. 

Edwin